OpenStudy (anonymous):

Solve: 4(x + 6) = 6(x − 2) plesase answer it , and if you can explain how you got that it would be great ..

OpenStudy (phi):

first use the distributive property multiply the 4 times each term inside the parens

OpenStudy (phi):

for example 4(x+6) means 4 times x plus 4 times 6 can you do that?

OpenStudy (anonymous):

4x6 = 24 ? like that ?

OpenStudy (phi):

you start with 4(x + 6) = 6(x − 2) now on the left side of the = sign do 4 times x plus 4*6 you get 4x+24= 6(x-2) can you do the right side (multiply 6 times each term inside the parens)

OpenStudy (anonymous):

6x2 = 12 ? 6times X = 6x ?

OpenStudy (phi):

what is 6*x?

OpenStudy (anonymous):

6x ?

OpenStudy (phi):

what is 6* -2?

OpenStudy (anonymous):

-12

OpenStudy (phi):

what is 6*(x-2) ?

OpenStudy (anonymous):

Not getting it ? 6times -2 = -12 and 6times X = 6X

OpenStudy (phi):

yes, but added together what is 6*(x-2) ?

OpenStudy (anonymous):

-12 - 6x ?

OpenStudy (phi):

where did you get -6x isn't 6*x just 6x what is 6*(x-2) ?

OpenStudy (anonymous):

I thought it was -12 minus 6x ?

OpenStudy (phi):

what is 6*(x-2) ?

OpenStudy (anonymous):

6x and -12

OpenStudy (phi):

people don't say and when they mean plus what is 6*(x-2) ?

OpenStudy (anonymous):

6x + -12

OpenStudy (phi):

ok, so we started with 4(x + 6) = 6(x − 2) we distribute the 4 (on the left) and the 6 on the right to get 4x+24= 6x + -12 next, add 12 to both sides of the equation.

OpenStudy (anonymous):

Where do i add the 12 ?

OpenStudy (anonymous):

to both sides of the equation

OpenStudy (phi):

4x+24= 6x + -12 add 12 to the left side and to the right side of the = sign

OpenStudy (anonymous):

so you will end up with 4x +36=6x

OpenStudy (anonymous):

then move your 4x to the other side and you get 36=6x-4x

OpenStudy (anonymous):

4x+36

OpenStudy (phi):

@maktaopvp write out the full equation, with the = sign 4x+24= 6x + -12 add 12 to the left side and to the right side of the = sign

OpenStudy (anonymous):

then add the -4x +6x and you get 36=2x then divide by 2 on each side

OpenStudy (anonymous):

4x+36=18x+-12 ?

OpenStudy (phi):

you need to add 12 to both sides where is the +12 on the right side?

OpenStudy (anonymous):

Okay i add 12 to the 24 and where else ?

OpenStudy (phi):

on the right side of the = sign

OpenStudy (anonymous):

so to the 6x ?

OpenStudy (phi):

yes, but write out the whole equation

OpenStudy (anonymous):

4x + 36 = 18 + -12 ??

OpenStudy (phi):

you start with 4x+24= 6x + -12 add +12 to both sides. NOTE: 6x+12 is not 18 and it is not 18x. just write +12 on both sides

OpenStudy (anonymous):

4x + 24 +12 = 6x + -12 +12 ??

OpenStudy (phi):

yes. now you can add the "pure numbers" (the terms without the x) what do you get? (Write out the whole equation)

OpenStudy (anonymous):

4x + 36= 6x and 12- + 12 = 0 ?

OpenStudy (phi):

yes, you get 4x + 36= 6x (because 12-12 = 0) now add -4x to both sides just write -4x on both sides

OpenStudy (anonymous):

4x + 36 + -4 = 6x + -4 ?

OpenStudy (phi):

no, not -4, -4x (the x is important)

OpenStudy (anonymous):

so its 4x + 36 + -4 = -4x ?

OpenStudy (anonymous):

its very easy 4x +24 = 6x-12 -2x=-36 x=18

OpenStudy (phi):

4x + 36= 6x write -4x on both sides

OpenStudy (anonymous):

4x + 36 -4x = 6x -4x ?

OpenStudy (phi):

yes now on the left side if you have 4 x's and you take away 4 x's, how many x's do you have?

OpenStudy (anonymous):

0

OpenStudy (phi):

so the left side is just 36 (because 4x-4x =0 and 36+0 is 36) 36= 6x-4x what is the right side?

OpenStudy (anonymous):

2x?

OpenStudy (phi):

yes, but write out the whole equation

OpenStudy (anonymous):

36= 2x

OpenStudy (phi):

OK, if 2 x's are 36 how much is 1 x?

OpenStudy (anonymous):

18

OpenStudy (phi):

x=18 is the answer. the way to do it is 36=2x divide both sides by 2 36/2 = 2x/2 18=x

OpenStudy (anonymous):

ahh okay thanks alot

OpenStudy (phi):

If you go over what we did, you might be able to solve other problems like this one.

OpenStudy (anonymous):

please dont be so hyper